OpenStudy (anonymous):

how to solve using change of variable: lim x->1 x^1/6 -1 / x^1/3 -1

OpenStudy (anonymous):

Let \[u = x^1/6\] then \[u^2 = x^1/3\] change limit u(1) = 1 lim u --> 1 u-1/(u-1)(u+1) \[\lim_{u \rightarrow 1} (1/u+1)=1/2\]